If 1.000 x 10^23 atoms of copper reacted with nitric acid, how many grams of water would be produced?
1
Expert's answer
2020-05-06T13:58:04-0400

"4HNO_3(l)+Cu(s) \rightarrow Cu(NO_3)_2(aq) + 2NO_2(g)+2H_2O(l)"

Convert number of atoms of copper to moles

"1.000\times 10^{23} (\frac{1 mole}{6.022\times10^{23}})=0.16606 moles"

Convert moles of copper to moles of water:


"0.16606 moles Cu (\frac{2 moles H_2O}{1 moles Cu})=0.33212 moles H_2O"


Convert moles of water to mass of water


"0.33212 moles H_2O (\frac{18.0.15 g H_2O}{1 mole H_2O})= 5.983 g H_2O"